OpenStudy (anonymous):

Two Awards and Being your fan ..... Is there any way 2 +2 = 5?

OpenStudy (ryguy):

haha no there isnt

OpenStudy (anonymous):

However, you can program your TI-83+ calculator to display to be true ;)

OpenStudy (anonymous):

Nope :(

OpenStudy (anonymous):

No haha

OpenStudy (anonymous):

Nossa,fiz este teste para testar seus conhecimentos e ninguém tem capacidade de responder....

OpenStudy (anonymous):

Can't understand you, but nope there isnt.

OpenStudy (anonymous):

Rir em voz alta :)

OpenStudy (anonymous):

I'm bad at Portuguese :P

OpenStudy (anonymous):

I suppose you could write up a stupid proof logically proving that it is true, but it will never be true and you cannot prove something incorrect.

OpenStudy (anonymous):

2+2 é igual a 5??? Vamos verificar: Começamos com a seguinte igualdade, que é verdadeira: 16-36 = 25-45 Somamos (81/4) nos dois lados, o que não altera a igualdade: 16-36+(81/4) = 25-45+(81/4) Isso pode ser escrito da seguinte forma: (trinômio quadrado perfeito) (4-(9/2))2 = (5-(9/2))2 Tirando a raiz quadrada em ambos os lados temos: 4-(9/2) = 5-(9/2) Somando (9/2) nos dois lados da igualdade temos: 4 = 5 Como 4=2+2 chegamos a seguinte conclusão: 2+2=5

OpenStudy (anonymous):

Can you type in English please? :)

OpenStudy (anonymous):

I mean I understand some of what your saying. I'm better at French though :)

OpenStudy (anonymous):

Français est génial, vous devriez essayer d'apprendre :)

OpenStudy (anonymous):

As I said. Stupid proofs.

OpenStudy (anonymous):

Sorry if I can not type anything in english,because I´m not in the country anymore ,I´m from Brazil

OpenStudy (anonymous):

Thing is, Assuming that (4-(9/2))2 = (5-(9/2))2 is actually (4-(9/2))^2 = (5-(9/2))^2 and that you applied the square root to both side, you would have gotten \[\pm (4-(9/2)) = \pm (5-(9/2))\]

OpenStudy (anonymous):

This is quite a stupid thread. There is NO way you can have 2 + 2 = 5 Like there's no way other than what @Grazes said, "Stupid Proofs" Can you close the question? @Jordannyah

OpenStudy (anonymous):

I replied and gives everything right to this type of account, just be smart ...